# custom.bib

### v2

@@ 4.1 Model and Training Parameters
In order to isolate the effect of changing the position encoding, we use a standard [^1] https://github.com/kentaroy47/vision-transformers-cifar10 transformer backbone modified to be able to switch between relative position encoding types. We use the standard backbone sizes of ViT-Tiny, ViT-B and ViT-L [dosovitskiy2020image]. All experiments use RandAugment [cubuk2020randaugment]. We avoid using pre-trained weights in order to maximize the comparability of results between methods. In order to ensure a fair comparison, we explicitly avoid tuning hyperparameters with LieRE and use the same default hyperparameters for all experiments (Appendix B).
4.2 Datasets and Tasks
Our experiments are designed to evaluate the efficacy of LieRE as a position encoding across both 2D and 3D data. We evaluate LieRE on the classification of 2D (images) and 3D (videos) data. For 3D data and ImageNet-1k (2D), we focus on accuracy. For CIFAR-100 (2D), where training is less resource intensive, we also evaluate LieRE’s data and training compute efficiency.
4.2.1 2D Classification
For 2D data we evaluate performance on the CIFAR-100 and ImageNet-1k image classification task [krizhevsky2009learning, deng2009imagenet]. We partition our evaluation of performance into four parts. In the first part, we examine accuracy across a range of model architectures on both CIFAR-100 and ImageNet-1k and compare to absolute [dosovitskiy2020image] position encoding and recent related work, RoPE-Mixed [heo2024rotary], VisionLlama embeddings [chu2024visionllama].
In the second part, we take advantage of the relatively modest amount of compute resources necessary to train a model for CIFAR-100 to examine LieREs impact on data efficiency. We also measure training compute efficiency by comparing the number of training steps necessary to achieve a fixed level of validation accuracy.
In the third part, we evaluate the impact of LieRE with various scales. We vary the capacity of the transformer backbone with corresponding to ViT-T, ViT-B and ViT-B, as proposed in [dosovitskiy2020image], the number of learned LieRE generators (one, per-attention-head and per-layer) and the LieRE generator capacity by imposing a block diagonal structure on the generator that allows us to vary the added capacity.
In the fourth part, we measure how much different models depend on the positional information in the image/video by shuffling the patches. A higher accuracy drop with randomly shuffled patches means the model relies more on the positions of the patches during inference.
4.2.2 3D Classification
In this section, we introduce Rotary Position Encodings for 3D data and compare LieRE-based transformers with transformers with RoPE-mixed and absolute encoding similar to the previous section [arnab2021vivit, heo2024rotary]. For the 3D experiments, we examine video classification performance in the UCF101 dataset [soomro2012ucf101]. Again, we did not optimize any hyperparameters for the LieRE model and used the dataloader from [tong2022videomae]. The full set of hyperparameters may be found in appendix B.
5 Results
5.1 Accuracy
For 2D images, we show that the LieRE-based transformer outperforms the DeiT by a relative performance improvement of 9.4% [touvron2022deit], RoPE adaptation in VisionLlama [chu2024visionllama] by 6.7% and RoPE-Mixed by 9.7% on CIFAR-100 [heo2024rotary].
For 3D inputs (video), we observe an relative accuracy improvement of LieRE-based transformer over the absolute position embedding of up to 25.5 % and RoPE-inspired position encodings at least to 6.1% in accuracy.
@@ 5.1 Accuracy
Table 1: 2D image and 3D video classification Top-1 Accuracy (95% confidence intervals) results. All models use 85.1M parameters for 2D tasks and 88.7M parameters for 3D task [krizhevsky2009learning, deng2009imagenet, soomro2012ucf101, rsna2019data] equivalent to DeiT, ∗∗ equivalent to Vivit (spatio-temporal).
@@
Abs. Pos. E.∗,∗∗ 63.9 (62.9-65.8) 66.1 (65.7-66.5) 42.7 (42.3-43.1)
VisionLlama RoPE 65.5 (64.6-66.5) 65.4 (65.0-65.8) 50.5 (50.1-50.9)
RoPE-Mixed 63.7 (61.8-65.6) 65.9 (65.5-66.3) 45.7 (45.2-46.0)
LieRE-Commute 66.9 (66.0 - 67.9) 68.4 (67.9-68.9) 52.1 (51.7-52.5)
LieRE 69.9 (68.9-70.8) 69.5 (69.1-69.9) 53.6 (53.2-54.0)
5.2 Data efficiency
We observe that transformers based on LieRE exhibit greater data efficiency compared to leading transformer architectures for 2D images on the CIFAR-100 dataset. For instance, when using 40% of the dataset, the LieRE-based transformer outperforms all comparable methods by 4.5 percentage points. Additionally, there is a mild trend indicating that the performance gap widens as the amount of data decreases (Figure 3(b)).
(a) Performance scales with generator dimension the number of LieRE parameters, compared to LieRE with one skew-symmetric matrix learned for the entire model (stem).
(b) Data ablation on for different position embeddings on CIFAR-100.
5.3 Model Scaling
We examine three axes of capacity scaling: transformer backbone parameters, LieRE generator parameters and learning separate LieRE generators across heads and layers. Here we find a surprising result that RoPE-mixed can be substantially improved by sharing parameters across attention heads and layers and not learning them separately.
5.3.1 Transformer backbone capacity
We also evaluate how the inclusion of LieRE affects performance across model sizes on the CIFAR-100 in Figure 4(a). We observe that LieRE retains a statistically significant lead in performance across all three model sizes.
@@ 5.3.1 Transformer backbone capacity
Refer to caption (a) Performance behavior on CIFAR-100 (2D Image Classification) over ViT-Tiny (22M), ViT-Base (85M), ViT-Large (302M) for LieRE Rope-Mixed and Absolute Encoding (Appendix, table 5).
Refer to caption (b) The LieRE spatial encoding allows the model to match the performance of absolute position encodings with substantially less training time.
5.3.2 Generator parameters
@@
LieRE adds a small amount of capacity to the model ($4032$ parameters for the ViT-B backbone we use for most experiments), leading to the natural question of how helpful the marginal capacity is.
We vary capacity by enforcing a block diagonal structure on the generator. varying the block size allows us to approach "full" LieRE. Recall that using $2\times 2$ blocks recovers exact commutativity and is referred to as LieRE-commute. This configuration substantially outperforms RoPE-Mixed, despite differing only in that it uses the same parameters across heads and layers and RoPE-Mixed learns separate parameters.
In figure 3(a) we evaluate accuracy versus block dimension. We observe the steepest increase at the very start, though accuracy continues to grow as the generator dimension increases.
5.3.3 Sharing LieRE Parameters Across Heads and Layers
We also examine whether learning separate positional encoding per attention head and per layer improves performance on CIFAR-100 in table 2. We conduct this experiment for both RoPE-Mixed and LieRE and observe that in both cases it is best to learn one set of parameters across heads and layers. This is the mechanism by which we are able to show larger improvements from adopting RoPE-Mixed than the original paper [heo2024rotary]. We name this transformed form of RoPE-Mixed LieRE-Commute, since it is equivalent to LieRE with generator dimension 2 (see Figure LABEL:fig:generator_dataset).
Table 2: Layer, Head and Stem Learnable Rotary Embeddings on CIFAR-100, ViT-B, equivalent to LieRE-Commute
@@ 5.3.3 Sharing LieRE Parameters Across Heads and Layers
FLOP Stem Heads Layers Rope-Mixed LieRE Reference
5.684G 63.0 64.5 [heo2024rotary]
5.684G 63.6 65.1
5.613G 68.3 69.9
5.613G 66.9 69.9 (ours)
5.4 Compute efficiency
Training transformers can necessitate substantial computational resources, which can hinder equitable access to research and development of machine learning methods. We demonstrate that the LieRE-based transformer requires 3.5 times less training epochs on Cifar100 to achieve comparable performance to the Absolute Position Embedding baseline (as used in DeiT III [touvron2022deit]). This represents the largest reduction in training time compared to recent works such as VisionLlama and RoPE-Mixed [chu2024visionllama, heo2024rotary]. We note that this is the first training efficiency comparison of these recent methods. Figure 4(b) shows the amount of allowable compute reduction to achieve the same accuracy achieved by absolute position encodings (DeiT baseline) after 200 epochs. LieRE demonstrates the largest win, allowing a 3.5X reduction in training compute to achieve the same accuracy.
5.5 FLOP analysis
Position Enc. ViT-Tiny (22M) ViT-Base (85M) ViT-Large (302M)
Abs. Pos. E. 0.963G 5.607G 19.856G
RoPE-Mixed 0.963G (+0.001%) 5.609G (+0.036%) 19.863G (+0.035%)
LieRE-Commute (ours) 0.963G (+0.000%) 5.608G (+0.018%) 19.858G (+0.005%)
LieRE (ours) 0.970G (+0.727%) 5.613G (+0.107%) 19.869G (+0.065%)
@@ 5.6 Patch shuffling
We evaluate models using the decline in accuracy when evaluating on shuffled patches. We observe the most significant decline LieRE-based transformers, leading to the conclusion that LieRE models are most capable at using positional information. The complete results are displayed for CIFAR-100 and table for UCF101 (Table 4) .
@@
Abs. Pos. E. 63.9 19.6 69.3 42.7 42.5 0.0
VisionLlama RoPE 65.8 29.7 54.8 50.5 36.2 28.5
RoPE-Mixed 63.6 38.7 39.0 45.7 37.9 17.1
LieRE-Commute 66.9 16.5 75.4 52.1 37.2 28.6
LieRE 69.9 15.9 77.1 53.6 24.7 53.9
@@ Appendix B Experiment Hyperparameters
The backbone for all experiments is configured as ViT-B, with 12 layers, a hidden dimension of 768, and an intermediate dimension of 3096. We use a dropout of 0.1. We used CLS pooling in our implementation to facilitate comparability with existing literature in the field. Further experiments revealed substantial performance improvement with mean pooling and LieRE. We use the pytorch lightning framework for all experiments [falcon2019pytorch].
@@
The CIFAR experiments where trained on 8xL4 GPUs with 24GB of VRAM each and all took under 30 minutes to complete. The generator scaling experiment was conducted using RTX6000 GPUs. The ImageNet experiments were trained on 8xL40 GPUs and all took less than 2 days and 5 hours of runtime including time lost due to preemption and resource sharing. We use a cosine learning rate schedule with an initial learning rate of $41E-$ 4 and train for 200 epochs. We use an effective batch size of $512$. We use a patch size of $4\times 4$ on the original $32\times 32$ image for CIFAR-100 and a patch size of $16\times 16$ on the randomly cropped and resized $224\times 224$ image. All vision experiments used RandAugment [cubuk2020randaugment]. We use the ADAM optimizer with betas of 0.9 and 0.999 and $\epsilon=1e-8$. The hyperparameters were tuned with RoPE-mixed and selected before conducting the LieRE trainers as to ensure a fair comparison.
@@
The 3D classification experiments were conducted on either $1008\times A$ 100 40GB GPUs or $1004\times A$ 100 80GB GPUs with the effective batch size held constant either by using a gradient accumulation or increasing the batch size. Similar to 2D classification, we use an initial learning rate of $41E-$ 4 with a cosine decay, trained for $200$ epochs, and had a total batch size of $64$ and a patch size of $2\times 16\times 16$ on the randomly cropped and resized $32\times 224\times 224$ video/image. We use the ADAM optimizer with betas of 0.9 and 0.999 and $\epsilon=1e-8$.
B.3 Scaling Model Size
Table 5: Comparison of Position Encoding Methods for Different ViT Models, Accuracy (bootstrapped 95%CI)
Position Encoding ViT-Tiny ViT-Base ViT-Large

### v3

@@ 5 Experiments
In order to isolate the effect of changing the position encoding, we use a plain standard transformer backbone and training regime [^1] https://github.com/kentaroy47/vision-transformers-cifar10 with comparable baseline performance on CIFAR100 and ImageNet with (Lee et al., 2022). We implement relative position encoding types including LieRE, Rope-Mixed, Visionllama, and absolute position encodings. We use the standard backbone sizes of ViT-Tiny, ViT-B and ViT-L (Dosovitskiy et al., 2020). All experiments use RandAugment (Cubuk et al., 2020). We avoid using pre-trained weights in order to maximize the comparability of results between methods. In order to ensure a fair comparison, we explicitly avoid tuning hyperparameters and use the same default hyperparameters for all experiments (Appendix B). We evaluate two versions of LieRE, distinguished by the basis matrix tile sizes of 64 and 8, referred to as LieRE64 and LieRE8, respectively. Notably, a tile size of 2 corresponds to RoPE-Mixed.
5.1 2D Classification
For 2D data we evaluate performance on the CIFAR-100 and ImageNet-1k image classification task (Krizhevsky et al., 2009; Deng et al., 2009). We further analyse different backbone models sizes and data ablation effects. We examine accuracy both CIFAR-100 and ImageNet-1k and compare to absolute (Dosovitskiy et al., 2020) position encoding and recent related work, RoPE-Mixed (Heo et al., 2024), VisionLlama embeddings (Chu et al., 2024). For 2D image classification tasks, we demonstrate that the LieRE-based transformer achieves a relative performance improvement in accuracy of 10.0% over standard absolute position encoding (Touvron et al., 2022), 7.3% over RoPE adaptation in VisionLlama (Chu et al., 2024), and 2.2% over RoPE-Mixed (Heo et al., 2024) on CIFAR-100, with similar accuracy trends observed on ImageNet (Deng et al., 2009) (table 1).
@@
Table 1: 2D image and 3D video classification Top-1 Accuracy (95% confidence intervals) results. All models use 85.2M parameters for 2D tasks and 88.7M parameters for 3D task (Krizhevsky et al., 2009; Deng et al., 2009; Soomro et al., 2012; Stein et al., 2019) equivalent to DeiT architecture, ∗∗ equivalent to Vivit (spatio-temporal) architecture.
@@
Abs. Pos. E.∗,∗∗ 63.9 (62.9-65.8) 66.1 (65.7-66.5) 40.9 (40.5-41.3)
VisionLlama RoPE 65.5 (64.6-66.5) 65.4 (65.0-65.8) 45.0 (44.6-45.4)
RoPE-Mixed 68.8 (67.9-69.7) 68.8 (68.4-69.2) 46.3 (45.9-46.7)
LieRE8 70.3 (69.4-71.2) 69.6 (69.2-70.0) 47.0 (46.6-47.4)
LieRE64 70.0 (69.1-70.9) 69.3 (68.9-69.7) 44.7 (44.3-45.1)
We vary the capacity of the transformer backbone with corresponding to ViT-Tiny, ViT-B and ViT-L, as proposed in (Dosovitskiy et al., 2020). Additionally, we analyze the impact of incorporating LieRE on performance across different model sizes on the CIFAR-100 dataset, as shown in table 5. Our results demonstrate that LieRE8 slightly outperforms alternatives across all evaluated model sizes.
@@
Figure 5: Performance behavior on CIFAR-100 (2D Image Classification) over ViT-Tiny (22M), ViT-Base (85M), ViT-Large (302M) for LieRE Rope-Mixed and Absolute Encoding (Appendix, table 5).
Our experiments reveal a pattern in how learnable relative position encodings perform under limited data conditions. Specifically, when we varied the available training data from 20% to 90% of the CIFAR-100 dataset, both LieRE and RoPE-Mixed demonstrated significantly better data efficiency compared to traditional transformer approaches for 2D image classification (figure 6).
Figure 6: Performance comparison of different positional embedding methods across varying dataset sizes. The plot shows the peak accuracy achieved by LieRE8, LieRE64, RoPE-Mixed, VisionLlaMA, and Absolute PE when trained on different fractions of the CIFAR-100 dataset. Both LieRE variants and RoPE-Mixed consistently outperform other methods, with their advantage becoming particularly pronounced in data-scarce scenarios.
5.2 Synthetic Spatial Classification
Figure 7: The model is asked to identify the direction of the arrow pointed to by the base of the Y. Here said arrow points down.
Recent observations indicate that even advanced models like ChatGPT-4 and Claude Sonnet 3.5 struggle with basic spatial reasoning tasks. To investigate whether these limitations stem from positional encoding mechanisms, we designed a synthetic image classification task (Shah et al., 2024). The task presents a $108\times 108$ pixel image containing a $9\times 9$ grid (81 cells). Within this grid, we randomly place eight arrows and six letters (A, B, C, D, E, and Y), ensuring that one arrow is placed in the direction of the base of the ”Y”. The objective is to identify the direction of this specific arrow. To introduce visual complexity, we include four spurious letters and seven additional arrows as distractors. Figure 7 illustrates an example of this task setup.
We verified that none of ChatGPT 4o, Claude Sonnet 3.5 and Gemeni Pro 1.5 are able to solve this task.
We train the models on 800,000 examples and observe that they generally converge after the first 400,000. The only exception to this is absolute position encodings, where we have variants trained on 800,000 and
Table 2: Comparing different positional encodings on synthetic task on base model (85M) equivalent to DeiT architecture across different resolutions, $108\times 108$, $168\times 168$, $276\times 276$
Method 108 168 276
Abs. Pos. E. 45.1 (43.1-47.7) 41.0 (39.0-43.2) 40.1 (38.2-42.0)
Abs. Pos. E. (2M examples) 47.2 (45.2-49.1) - -
VisionLlama RoPE 46.4 (44.0-47.9) 48.6 (46.0-49.9) 48.6 (46.2-50.1)
RoPE-Mixed 100 (99.5-100) 98.6 (97.6-98.7) 88.6 (87.4-89.9)
LieRE8 99.5 (99.2-99.8) 99.7 (99.4-99.9) 99.7 (99.5-99.9)
LieRE64 100 (99.5-100) 100 (99.6-100) 100 (99.7-100)
Table 2 outlines the performance of different positional encoding methods on a synthetic task using a base model (85M parameters) equivalent to the DeiT architecture. We evaluate the models across three different input resolutions ($108\times 108$, $168\times 168$, and $276\times 276$ pixels), revealing differences in scalability and effectiveness. While absolute positional encoding shows degraded performance as resolution increases (from 45.1% to 40.1%) and RoPE-Mixed demonstrates strong but degrading performance at higher resolutions (from 100% to 88.6%), both LieRE variants maintain near-perfect accuracy across all resolution scales, with LieRE64 achieving 100% accuracy consistently, suggesting superior resolution invariance of our Lie group-based approach. Qualitative analysis shows that absolute and visionllama position encoding attend less clearly to the ”Y” token than Rope-Mixed and LieRE. Please refer to the appendix for attention map examples, Figure 15 and Figure 15.
5.3 3D Classification
In this section, we introduce Rotary Position Encodings for 3D data and compare LieRE-based transformers with transformers with RoPE-Mixed and absolute encoding similar to the previous section (Arnab et al., 2021; Heo et al., 2024). For the 3D experiments, we examine video classification performance in the UCF101 dataset (Soomro et al., 2012). Again, we did not optimize any hyperparameters for the LieRE model and used the dataloader from (Tong et al., 2022). The full set of hyperparameters may be found in appendix B. We observe a relative accuracy improvement of the LieRE-based transformer of up to 15.1% compared to absolute position embeddings and at least 1.5% compared to RoPE-inspired position encodings (table 1).
5.4 LieRE parameter
@@
LieRE adds a small amount of capacity to the model ($580k$ parameters 0.68% for the ViT-B backbone we use for most experiments), leading to the natural question of how helpful the marginal capacity is. To answer this question, we vary the density of the learned basis and the learnability across heads and layers.
We vary capacity by enforcing a block diagonal structure on the skew-symmetric basis. Varying the block size allows us to approach LieRE64. Recall that using $2\times 2$ blocks recovers exact commutativity and is identical to RoPE-Mixed. As RoPE-Mixed is sensitive to the initialization of the weights, we initialize all generator weights mirroring the original presentation of RoPE-Mixed (Heo et al., 2024). In figure 8 we evaluate accuracy versus block dimension (Appendix Table 8).
Figure 8: Performance varies with skew-symmetric basis learnable dimensions, referred to as LieRE parameters. This is equivalent to increasing the tile size in the skew-symmetric basis ($2\times 2,4\times 4,8\times 8,16\times 16,32\times 32,48\times 48$). For both 2D (CIFAR-100) and 3D (UCF101) LieRE with tile size $8\times 8$ performs superior.
Table 3: Accuracy with parameter sharing over heads and layers for ViT-B sized models on CIFAR-100.
@@
FLOP All Shared Shared Across Layers Shared Across Heads RoPE-Mixed LieRE64 LieRE8
5.684G 68.8 70.0 70.3
5.684G 68.7 69.5 69.8
5.613G 69.5 69.7 69.7
5.613G 68.3 69.4 69.5
Table 3 presents the impact of sharing LieRE parameters across attention heads and layers on CIFAR-100 classification performance for RoPE-Mixed, LieRE64 and LieRE8 . We evaluate whether learning separate positional encodings for each attention head and layer provides performance benefits. We observe that learning across heads and layers yields superior performance followed by learning across layers.
@@ 5.5 Patch shuffling
We evaluate models using the decline in accuracy when evaluating on shuffled patches. We observe the most significant decline LieRE-based transformers, leading to the conclusion that LieRE models are more capable at using positional information. The complete results are displayed for CIFAR-100 and table for UCF101 (table 4).
@@
Abs. Pos. E. 63.9 19.6 69.3 40.9 39.5 0.0
VisionLlama RoPE 65.5 29.7 54.8 45.0 37.0 17.7
RoPE-Mixed 68.8 17.1 75.1 46.3 28.2 39.1
LieRE8 70.3 12.3 82.5 47.0 27.8 40.9
LieRE64 70.0 10.8 84.6 44.7 28.0 37.4
Figure 9: Validation accuracy comparison between LieRE 64 and RoPE-Mixed positional embeddings across different image resolutions on ImageNet. Left: Models trained at $224\times 224$ resolution. Right: Models pre-trained at $224\times 224$ and fine-tuned at $256\times 256$ resolution. Both approaches show similar performance trends up to $320\times 320$, after which LieRE 64 demonstrates significantly better accuracy retention at higher resolutions, particularly in the fine-tuned scenario.
5.6 Multi-resolution Classification
In this section we compare the ability of methods to generalize to image resolutions not seen during training. We evaluate two training recipes inspired by (Heo et al., 2024). The first recipe matches the rest of the paper and consists of training the models on images of size $224\times 224$ for 200 epochs. The second adds an additional fine-tuning step at size $256\times 256$ for 30 epochs. The full details can be found in appendix B. We evaluate the accuracy on the ImageNet validation set with varying inference resolutions. Specifically, we scale the input images to resolutions of $196\times 196$, $256\times 256$, $320\times 320$, $384\times 384$, and $448\times 448$ pixels per dimension, and present the resulting accuracies in figure 9. For position assignment, we adopt a sequential approach where token positions are scaled proportionally to the image dimensions. For example, doubling the length of an image in each dimension doubles the range of positional indices. This method outperforms rescaling positions to a fixed range, as demonstrated by superior results for both RoPE-Mixed and LieRE across the evaluated training recipes.
6 Conclusion
@@ Appendix B Experiment Hyperparameters
The backbone for all experiments is configured as ViT-B, with 12 layers, a hidden dimension of 768, and an intermediate dimension of 3096. We use a dropout of 0.1. We used CLS pooling in our implementation to facilitate comparability with existing literature in the field. Further experiments revealed substantial performance improvement with mean pooling and LieRE. We use the pytorch lightning framework for all experiments (Falcon, 2019).
@@
The CIFAR experiments where trained on 8xL4 GPUs with 24GB of VRAM each and all took under 30 minutes to complete. The basis capacity scaling experiment was conducted using RTX6000 GPUs. The ImageNet experiments were trained on 8xL40 GPUs and all took less than 2 days and 5 hours of runtime including time lost due to preemption and resource sharing. We use a cosine learning rate schedule with an initial learning rate of $41E-$ 4 and train for 200 epochs. We use an effective batch size of $512$. We use a patch size of $4\times 4$ on the original $32\times 32$ image for CIFAR-100 and a patch size of $16\times 16$ on the randomly cropped and resized $224\times 224$ image. All vision experiments used RandAugment (Cubuk et al., 2020). We use the ADAM optimizer with betas of 0.9 and 0.999 and $\epsilon=1e-8$. The hyperparameters were tuned with RoPE-Mixed and selected before conducting the LieRE trainers as to ensure a fair comparison.
@@
The 3D classification experiments were conducted on either $1008\times A$ 100 40GB GPUs or $1004\times A$ 100 80GB GPUs with the effective batch size held constant either by using a gradient accumulation or increasing the batch size. Similar to 2D classification, we use an initial learning rate of $41E-$ 4 with a cosine decay, trained for $200$ epochs, and had a total batch size of $64$ and a patch size of $2\times 16\times 16$ on the randomly cropped and resized $8\times 224\times 224$ video/image. We use the ADAM optimizer with betas of 0.9 and 0.999 and $\epsilon=1e-8$.
B.3 Multi-resolution Classification
The second training recipe consists of 30 epochs with an initial learning rate of 1E-5 with a cosine decay. This mirrors the DEIT III training reciple that first pretrains at a lower resolution and finetunes at a higher resolution.
B.4 CIFAR-100 Performance Across Model Scales
We also evaluate how the inclusion of LieRE affects performance across model sizes on the CIFAR-100 in table 5. We observe that LieRE retains a statistically significant lead in performance across all three model sizes.
B.4.1 Basis parameters
Table 5: Comparison of Position Encoding Methods for Different ViT Models Sizes on CIFAR-100, Accuracy (bootstrapped 95%CI)
Position Enc. ViT-Tiny (22M) ViT-Base (86M) ViT-Large (302M)